OpenStudy (anonymous):

Write the expression 2^3/5 as a radical.

OpenStudy (anonymous):

Simplify the expression . 

Nnesha (nnesha):

hint: exponent rule \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]to get rid of negative exponent change that to its reciprocal when yo do that sign of exponent would change 2nd: exponent rule \[\huge\rm \frac{ x^m }{ x^n } =x^{m-n}\]

OpenStudy (haseeb96):

Boht ache @Nnesha

Nnesha (nnesha):

shukria o^_^o

Nnesha (nnesha):

btw for first question \[\huge\rm \sqrt[n]{x^m}=x^\frac{ m }{ n }\] number at denominator would be root # numerator (m) is an exponent of variable or number which is under the root

OpenStudy (mathmate):

@jessica_the_meow According to the rule of precedence PEMDAS, 2^3/5 means \(\large\frac{2^3}{5}\) To write \(\large2^{\frac{3}{5}}\) correctly, parentheses are required, as in 2^(3//5). Even though other tutors understand this time what you meant by the context, omission of parentheses can get you in trouble by solving the wrong question.